### v1

# Laser-driven Ultrafast Dynamics of a Fractional Quantum Hall System

Abstract
Fractional quantum Hall (FQH) systems are strongly interacting electron systems with topological order. These systems are characterized by novel ground states, fractionally charged and neutral excitations. The neutral excitations are dominated by a low-energy collective magnetoroton mode. Here we derive and use a quasi-one-dimensional model to investigate the ultrafast nonequilibrium dynamics of a laser-driven FQH system within a two-Landau-level approximation. As opposed to the traditional and synthetic bilayers, our model accounts for interactions where electrons can scatter from one Landau-level to another. By performing exact time evolution of the system, we create an out-of-equilibrium state following the laser pulse that shows rich physics. Our calculations show the presence of non-trivial excited modes. One of these modes is electromagnetically active and represent density oscillations of magnetoplasmon mode. Another mode is identified by evaluating the overlap of the initial state and the out-of-equilibrium state following the laser pulse with a quadrupole operator. This mode is analogous to the chiral-graviton mode for FQH systems recently measured in experiments [Nature 628, 78 (2024)]. Our results show that a linearly-polarized pulse field can excite the graviton mode when inter-Landau level scattering occurs.
Introduction.— Fractional quantum Hall (FQH) systems are a prototypical example of strongly correlated physical systems with rich novel phenomena like topological order, quantum geometry and excitations that follow fractional statistics Haldane (2011); Arovas et al. (1984). The neutral excitations of FQH systems are dominated by a collective coherent oscillation mode usually called Girvin-MacDonald-Platzman magnetoroton mode Girvin et al. (1986, 1985). In the long-wavelength limit $k\to 0$, the magnetoroton becomes a quadrupole excitation that carries $L=2$ angular momentum along with the excitation of an underlying quantum geometry metric analogous to the passing of gravitational waves in a medium Haldane (2011). Another novel consequence of the presence of quantum geometry and topology in FQH systems is the quantization of Hall viscosity Haldane (2009); Pu et al. (2020); Read and Rezayi (2011). FQH phase is characterized by the quantization of transverse component of resistivity at low temperatures. This is measured in 2D electrons with density $\rho$ in the presence of high perpendicular magnetic field $B$ at fractional filling $\nu=\rho h/eB$ ($h$ being Planck’s constant and $e$ is the electron charge). Among many odd-denominator fractional filling states observed experimentally Tsui et al. (1982); Stormer et al. (1983) and theoretically explained Laughlin (1983); Haldane (1983), the first observed FQH state at $\nu=\frac{1}{3}$ is of particular interest. It was shown that fractional $\frac{1}{3}$ state can host charged excitations like quasi-particle (quasi-hole) of $e/3$ sometimes called anyons Arovas et al. (1984). It was not until recently that experiments performed on this FQH state gave first direct observation of exchange statistics of fractionally charged particles (anyons) Nakamura et al. (2020). Furthermore, interest in this state is rekindled due to the observation of two $e/3$ quasi-particles scattering into an electron and “Andreev reflected” quasi-hole Glidic et al. (2023). Besides its experimental accessibility, this state has also been readily studied theoretically on different geometries Yoshioka et al. (1983); Rezayi and Haldane (1994); Nakamura et al. (2012); Soulé and Jolicoeur (2012). Moreover, a long wavelength low-energy excitation mode of $\frac{1}{3}$ FQH state has been excited in inelastic linearly-polarized light scattering experiments Pinczuk et al. (1993). This mode was identified later on as a chiral-graviton mode that can be excited through circularly polarized light Liou et al. (2019). Another experiment has also excited this mode through two-photon processes in FQH-1/3 state Liang et al. (2024). Despite such experiments, the theoretical understanding of the dynamics of a multi-Landau-level FQH system has been limited.
In this Letter, we present ultra-fast dynamics of the FQH system is driven out-of-equilibrium by a laser pulse field. The system is initially in $\nu=\frac{1}{3}$ FQH state. We show that when all inter-two-level scattering terms are considered, a short time linearly-polarized light with frequency near Landau level gap can excite chiral-graviton mode below magnetoplasmon mode. We note that while optically driven synthetic bilayers have been recently proposed Ghazaryan et al. (2017); Cian et al. (2020) to host new types of FQH states Halperin (1983); He et al. (1993), as in conventional semiconductor bilayers Eisenstein et al. (1992); Suen et al. (1992); Bishop et al. (2007); Padmanabhan et al. (2009), the electronic systems are assumed to thermalize to the Floquet ground state in the rotating framework of the optical drive field. Therefore, the problem studied in the present work is quite distinct from the situations studied before.
Model Hamiltonian.— We consider a 2D spinless electronic system defined on the surface of torus with lengths $L_{x}$ and $L_{y}$, as shown in Fig. 1 in the presence of both a high perpendicular static magnetic field and a time-dependent electric field from the laser pulse. In the Landau gauge, both magnetic and electric fields can be described by the vector potential $\mathbf{A}=\mathbf{A}_{\rm static}(\mathbf{r})+\mathbf{A}_{\rm EM}(t)$ with $\mathbf{A}_{\rm static}=(0,Bx,0)$ and $\mathbf{A}_{\rm EM}=(0,A_{\rm EM}(t),0)$. The single-particle states in this gauge compatible with torus boundary conditions are given by Yang et al. (2012)
$\phi_{n,j}(\mathbf{r})=\frac{1}{\sqrt{{\cal N}_{n}}}\sum_{k}$ $e^{i\frac{(X_{j}+kL_{x})y}{\ell_{B}^{2}}}e^{-\frac{(x+X_{j}+kL_{x})^{2}}{2\ell_{B}^{2}}}$ (1)
$H_{n}\bigg{(}\frac{x+X_{j}+kL_{x}}{\ell_{B}}\bigg{)}$
Where $X_{j}=\frac{2\pi j}{L_{y}}\ell_{B}^{2}$ with $\ell_{B}=\sqrt{\hbar/eB}$ being the magnetic length. ${\cal N}_{n}=L_{y}\sqrt{\pi}2^{n}n!\ell_{B}$ is the normalization constant and $H_{n}$ is Hermite polynomial of order $n$. For each Landau level $n$, index $j$ runs from $10,1,...,N_{\phi}-$ 1, where $N_{\phi}=L_{x}L_{y}/(2\pi\ell_{B}^{2})$ is the number of flux quanta (Landau orbitals) passing through the system. Within this Landau orbital basis, we are able to write the system Hamiltonian as $\hat{H}=\hat{H}_{sp}+\hat{H}_{int}$, with
$\hat{H}_{sp}$ $=\omega_{c}\sum_{j}\hat{n}_{1,j}+\Gamma_{\rm EM}(t)[\sum_{j}c_{0,j}^{\dagger}c_{1,j}+\mathrm{H.c.}]\;,$ (2)
describing the coupling of electrons with the laser pulse field up to the first Landau level (LL1), and
$\hat{H}_{int}=\sum_{\{n_{i},j_{i}\}}V^{n_{1},n_{2};n_{3},n_{4}}_{j_{1},j_{2};j_{3},j_{4}}c_{n_{1},j_{1}}^{\dagger}c_{n_{2},j_{2}}^{\dagger}c_{n_{3},j_{3}}c_{n_{4},j_{4}}\;,$ (3)
describing the Coulomb interaction with $V_{n_{1},n_{2};n_{2},n_{4}}^{j_{1},j_{2};j_{3},j_{4}}$ the amplitude of two-particle scattering processes due to the repulsive interactions. Here the operators $c_{m,j}$, $c_{m,j}^{\dagger}$ ($\hat{n}_{m,j}{\equiv}c_{m,j}^{\dagger}c_{m,j}$) destroy or create an electron in $m$-th Landau level (LL) at $j$-th orbital localized around $2\pi j\ell_{B}^{2}/L_{y}$ along the axis of the cylinder. The quantity $\omega_{c}$ is the level-spacing between Landau levels 0-1 (see Fig. 1). In each Landau level, cyclotron frequency of electron is quantized in units of $\omega_{c}=eB/m^{*}$ ($m^{*}$ is electron’s effective-mass). The laser field coupling strength $\Gamma_{\rm EM}=\frac{eA_{EM}(t)}{\sqrt{2}m^{*}\ell_{B}}$. We also note for electrons interacting through Coulomb repulsion in the system, the Coulomb energy $E_{c}=e^{2}/\varepsilon\ell_{B}$ ($\varepsilon$ being di-electric constant) is also an important energy scale. The interaction part of our Hamiltonian Eq. (3) contains both intra- and inter-level scattering processes. Throughout the work, we use a short pulse form $A_{\rm EM}(t)=A_{0}e^{-\frac{(t-t_{c})^{2}}{2t_{d}^{2}}}\cos\big{[}\omega_{0}(t-t_{c})\big{]}$ where $t_{d}$ is time delay, $t_{c}$ is the center time for the pulse and $\omega_{0}$ is the frequency of incoming laser Fauseweh and Zhu (2020).
Figure 1: (a) Torus with lengths $L_{x}$ and $L_{y}$. Magnetic field is pointing along $\hat{z}$. (b) A schematic of a two-level FQH system. Initially, the system is in fractional $\nu=1/3$ state occupying the LLL ($n=0$). This state is then laser-driven creating excitations to the first Landau level ($n=1$). The energy difference between the two levels is $\omega_{c}$.
In the absence of the pulse field and for large positive $\omega_{c}$, our fermions are in the ground state of the system, containing $N$ electrons, such that the filling factor is $\nu{=}N/N_{\phi}{=}1/3$. Also the system Hamiltonian commutates with the center of mass-momentum operator $\hat{K}=\sum_{j,n}jc_{n,j}^{\dagger}c_{n,j}$ (in Landau gauge) and momentum transfer by the laser pulse is considered to be negligible. Hence, our many-body dynamics is contained in the momentum eigen-sector of the initial state.
Figure 2: Two-electron virtual processes that do not conserve Landau level index.
To include the effects of the out-of-equilibrium physics under the influence of a laser-pulse, the Hamiltonian includes both scattering processes that conserve Landau level index $n_{1}+n_{2}=n_{3}+n_{4}$ and ones that break the level index i.e. $n_{1}+n_{2}\neq n_{3}+n_{4}$. The latter is schematically illustrated in Fig. 2. This is the consequence of an electron scattering from one level to another and is ignored in traditional and synthetic bi-layers Ghazaryan et al. (2017). We avoid large scale Hilbert space by truncating Hamiltonian to LL1 and by considering low-intensity radiation. This assumption is further validated by another resonant inelastic light experiment for FQH $\frac{1}{3}$ state showing most features related to excitations from LLL to LL1 Pinczuk et al. (1993). Coulumb’s interaction potential for torus boundary conditions is $V(\mathbf{r})=\sum_{s,t}e^{2}/\varepsilon|\mathbf{r}+sL_{x}\hat{x}+tL_{y}\hat{y}|$ where $\hat{x}$, $\hat{y}$ are the units vectors along the two axes of the torus Yoshioka et al. (1983). We obtain the matrix elements
$V_{j_{1},j_{2},j_{3},j_{4}}^{n_{1},n_{2},n_{3},n_{4}}=\frac{\delta_{j_{1}+j_{2},j_{3}+j_{4}}^{{}^{\prime}}}{2L_{x}L_{y}}\sum_{\mathbf{q},\mathbf{q}\neq 0}\delta_{q_{y},t\frac{2\pi}{b}}\delta_{q_{x},s\frac{2\pi}{a}}\delta_{j_{1}-j_{4},t}^{{}^{\prime}}\frac{2\pi e^{2}}{\varepsilon q}C_{n_{1},n_{4}}(-\mathbf{q})C_{n_{2},n_{3}}(\mathbf{q})e^{-i2\pi s\frac{j_{1}-j_{3}}{N_{\phi}}}e^{-\frac{q_{x}^{2}+q_{y}^{2}}{2}\ell_{B}^{2}}$ (4)
Where $C_{0,0}(\mathbf{q})=1$, $C_{1,1}(\mathbf{q})=1-\frac{|q|^{2}\ell_{B}^{2}}{2}$ and $C_{1,0}(\mathbf{q})=\frac{q_{y}+iq_{x}}{\sqrt{2}}\ell_{B}=\big{[}C_{0,1}(-\mathbf{q})\big{]}^{*}$. Prime over a delta function indicates that it is defined modulo $N_{\phi}$. Fractional $1/3$ state Laughlin (1983) is the ground state of the Hamiltonian $\sum_{j_{i}}V_{j_{1},j_{2},j_{3},j_{4}}^{0,0,0,0}c_{j_{1}}^{\dagger}c_{j_{2}}^{\dagger}c_{j_{3}}c_{j_{4}}$ and is adiabatically connected to period-three charge-density $|100,100,100,100,...\rangle$ Seidel et al. (2005); Trugman and Kivelson (1985) (for short-range potential $\nabla^{2}\delta(\mathbf{r})$). The physics of FQH systems are usually studied in the limit of large level-spacing and the Hamiltonian for single Landau level is diagonalized. This is due the assumption that Coulomb interaction is weak to scatter electron from one level to another in the large magnetic field limit which is not always true Papić and Balram (2022). We have first verified that the ground state of of the system Hamiltonian involving virtual process for various values of level spacing is indeed the fractional-$\frac{1}{3}$ quantum Hall state (See Supplemental Material (SM) sup for more details).
[…]
$\hat{J}_{x/y}=\frac{\hbar(i/1)}{m\sqrt{2}}\sum_{j}\bigg{(}c_{1,j}^{\dagger}c_{0,j}\mp c_{0,j}^{\dagger}c_{1,j}\bigg{)}\;.$ (5)
The details of this quantity is given in the SM sup .
Figure 3: Color online. Dynamical observables for various laser frequencies $\omega_{0}$ (a) Number of excited electrons for various in-coming light frequencies in-set of (a) gives the form of $\Gamma_{\rm EM}(t)$ (b) Fidelity $|\langle\Psi(0)|\Psi(t)\rangle|$. (c) The y-component of current from Eq. (5). d) $x$-component of the current. The quantities are obtained for $N=6$ on torus with $L_{2}=10.63\ell_{B}$. The intensity is set to $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/E_{c}$. The units of time is the inverse of $E_{c}$. The usual values of $E_{c}$ is around $14\;{\rm meV}$ Pinczuk et al. (1993) which gives $47$ femto-seconds as the unit of time.
Figure 4: Spectrum of $d[\hat{J}_{y}(t)]/dt$ following the pulse with parameters $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/Ec$. The in-coming laser frequency is swept from $\omega_{0}=0.6-0.8\omega_{c}$. The behaviour of the current shows the excitation of a mode in addition to magnetoplasmon at frequency $w_{c}$ that is blue-shifted from $w_{c}$. The system size is set to $N=6$ and $L_{y}=\sqrt{2\pi 18}\ell_{B}$. Inset gives the spectrum for near-TT limit ($L_{y}=2\pi\ell_{B}$).
Results. We consider level-spacing $\omega_{c}<E_{c}$, i.e. the level spacing is less than Coulomb energy scale. The ratio of these energies $\kappa\equiv\frac{e^{2}}{\hbar\omega_{c}\epsilon\ell_{B}}$ is magnetic field dependent ranging from $2.6/\sqrt{B[T]}$ for Ga-As to $22.5/\sqrt{B[T]}$ for Al-As justifying the our approximation for fractional $\nu=1/3$ Hall system usually observed at $B\sim 10$ Tesla Sodemann and MacDonald (2013). In Fig. 3, we present observables like fidelity, mean-excitation number and spectrum of $d[\hat{J}(t)]/dt$ after the pulse for $\omega_{c}=E_{c}/2$ for $L_{y}=\sqrt{2\pi 18}\ell_{B}$. Fig. 3 (a-b) gives the excitations to LL1 and the fidelity. The current $\hat{J}_{x,y}$ in the Fig. 3 (c-d) remains oscillating with multiple periods after the pulse is finished. Fig. 4 gives the spectrum of current of ${\mathcal{J}}_{y}(\omega)={\mathcal{F}}_{t}\big{[}\frac{d}{dt}\hat{J}_{y}(t)\big{]}(\omega)$ for $L_{y}=\sqrt{2\pi 18}\ell_{B}$ (liquid limit) and $L_{y}=2\pi\ell_{B}$ (near-TT limit). There are two modes as seen from the spectrum for low-intensity short time pulses one at $\omega_{c}$ and another blue-shifted from $\omega_{c}$. The presence of these modes are robust against slight change in the incoming frequencies although their strength can change. The peak at the level-spacing (cyclotron-energy) $\omega_{c}$ is the plasmon equivalent of FQH state identified as $q\to 0$ limit of magnetoplasmon mode. This mode is due to single particle excitations from LLL to LL1 Pinczuk et al. (1993). In the experiments, another peak that is blue shifted from magnetoplasmon has also been observed Pinczuk et al. (1993). This mode was sometimes attributed to $q=0$ spin-flip inter-Landau-level excitation involving enhanced exchanged in spin-polarized state. We find blue-shifted peaks in our calculations are due to excitations of additional electron to $LL1$ although complete understanding is missing due to the presence of finite-size effects and the truncation to LL1.
Figure 5: Color online. Structure factor ${\mathcal{S}}(q_{x}=2\pi/L_{x},\omega)$ for different system sizes. The torus dimensions is taken $L_{x}=L_{y}=\sqrt{6\pi N}\ell_{B}$. The laser-frequency is $\omega_{0}=0.7\omega_{c}$ and other laser parameters are set as in Fig. 3.
The spectrum of structure factor ${\mathcal{S}}(\mathbf{q},\omega)$ is shown in Fig. 5. It shows the existence of low-energy modes for different system sizes. The peak at $\omega/\omega_{c}=1$ is magnetoplasmon mode discussed in the laser-field induced current spectrum. Significantly, there emerge new modes located at the energy around $0.2\omega_{c}$. We have also verified that in the low intensity regime of the laser pulse field, the magnetoplasmon mode and the new mode follow the linear and quadratic dependence of the field strength, respectively. To determine the nature of theses new low-energy modes in $S_{\mathbf{q}}(\omega)$, we evaluate the overlap $|\langle{\Psi(t)}|n\rangle|$ to analyze the weight of low-lying eigen-states $|n\rangle$ contained in the photo-induced excited state $|\Psi(t)\rangle$. The results are given in Table 1. We have found that low-energy excitations are dominated by quadrupole FQH graviton states. These states are identified by matrix-element $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$, where $\hat{O}^{(2)}$ is the operator with quadrupole structure and $|\Psi_{0}\rangle$ is our initial ground state before the application of laser pulse. The specific form of $\hat{O}^{(2)}$ is given as Liou et al. (2019)
$\hat{O}^{(2)}=\sum_{\{j_{i}\}}\delta_{j_{1}+j_{2},j_{3}+j_{4}}^{{}^{\prime}}\hat{O}^{j_{1},j_{2},j_{3},j_{4}}_{qd}c_{0,j_{1}}^{\dagger}c_{0,j_{2}}^{\dagger}c_{0,j_{3}}c_{0,j_{4}}\quad$ (6)
$\hat{O}^{\{j_{i}\}}_{qd}=\sum_{q=(s\frac{2\pi}{L_{x}},t\frac{2\pi}{L_{y}})}^{{}^{\prime}}\delta_{j_{1}-j_{4},t}^{{}^{\prime}}\bigg{(}q_{x}^{2}-q_{y}^{2}\bigg{)}\frac{2\pi}{q}e^{-\frac{q^{2}}{2}}e^{-i2\pi s\frac{j_{1}-j_{3}}{N_{\phi}}}$
Figure 6: Color online. (a) $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$ for near-TT limit $L_{y}=2\pi\ell_{B}$. Where $|\Psi_{0}\rangle$ is the FQH-1/3 state and $|n\rangle$ is the $n$-th excited state of Hamiltonian in Eq. (3). The inset of (a) gives the transition matrix element between FQH-1/3 state and the eigen-states of LLL. (b) Same as (a) but for quantum-liquid limit ($L_{y}=10.63\ell_{B}$) for square torus. The inset of (b) gives the matrix element for the liquid-limit in LLL. The system size is $N=6$ particles and the level-spacing is set to $0.5E_{c}$. Grey vertical line on plots is at energy $0.15E_{c}$.
In Fig. 6, we present the values of transition matrix element $\langle n|\hat{O}^{(2)}|\Psi_{0}\rangle$ (un-normalized) between FQH 1/3 state and $|n\rangle$ is the $n$-th excited state for two cases i) When $|n\rangle$ is the eigen-state of the system Hamiltonian involving all scattering processes and ii) when $|n\rangle$ is the eigen-state of LLL Hamiltonian where all scattering processes are confined to LLL. This transition matrix element represents the system transition rate due to an oscillating metric tensor analogous to the passing of a gravitation wave and hence usually termed as graviton Yang (2016). This graviton mode occurs for both near-TT and liquid limit as shown in Fig.  6 (a) and (b). It is robust against the system size (See SM  sup for details). Moreover, as seen from Fig. 6, the energy of graviton states $E_{grav}$, when virtual processes are included, is slightly lower than the graviton states of a system when all interaction processes are confined to LLL. Furthermore, we follow Ref. Liou et al., 2019 to analyze the handedness of this graviton and find it has the structure of $(q_{x}-iq_{y})^{2}$, suggesting its chiral nature (See SM  sup for the detail of analysis.). Previously, it has also been proposed  Kirmani et al. (2022); Liu et al. (2018) that graviton can be excited via geometric quench Liu et al. (2021). A variant of such graviton mode called chiral graviton mode has been observed in experiments recently for fractional Hall states at filling fractional $\frac{1}{3}$ and $\frac{2}{5}$ fillings Liang et al. (2024). Chiral graviton modes in FQH system has been previously excited by tuning the laser-frequency near resonance conditions between valence and conduction band in 2D quantum wells. Our work has demonstrated that chiral graviton modes can also be exited when the a short-time laser is tuned near Landau level gap which is easily accessible experimentally. Furthermore, the excitation of graviton mode can be observed along with magneto-plasmon mode when all scattering processes are considered. We suggest to use a circularly polarized probe light to experimentally detect the chirality of the graviton mode.
Table 1: The contribution of states in $|\Psi(t)\rangle$ after the pulse for $N=5$ and $N=6$. $\Delta E$ denotes the energy difference from the ground state. Overlap $|\langle n|\Psi(t)\rangle|$ gives the contribution of individual eigen-state $|n\rangle$ of the system Hamiltonian with energy $E_{n}$. $\omega_{c}=0.5E_{c}$.
Size  [^1] The parameters of laser is set to $(eA_{0}/\sqrt{2}m^{*}\ell_{B},\omega_{0})=(0.2E_{c},0.7\omega_{c})$ with $t_{d}=10E_{c}$ and $L_{y}=\sqrt{2\pi N_{\phi}}\ell_{B}$. Similar results have been found for $N=7$ and are also observed for near-TT limit. $E_{n}$ ($E_{c}$) Overlap $\Delta E$ $(E_{c})$ Nature  [^2] Due to large Hilbert space size, we do not have eigen-state corresponding to Blue-shifted peak for $N=6$. Similarly, for $N=7$ we only have low-energy quadrupole states.
$N=6$1.38 0.84 0.0 G.S.
- -1.296 0.092 0.084 Quadrupole
- -1.2828 0.05 0.097 Quadrupole
- -0.876 0.4288 $w_{c}$ Plasmon
N=5 -1.067 0.882 0.0 G.S.
- -0.9377 0.0934 0.129 Quadrupole
- -0.562 0.396 $0.504\approx\omega_{c}$ Plasmon
- -0.091 0.086 0.543 Blue-shifted
Conclusions.— Through exact-diagonalization, the laser-induced dynamics of FQH state at $\nu=1/3$ has been presented. We showed that laser-induced excitations of FQH states have rich physics. On the one hand, some excitations are coupled to electromagnetic field and can cause emission like the plasmon mode. On the other hand, we have shown that when scattering from LLL to LL1 is included in interactions, it is possible to excite neutral-modes of LLL through linearly polarized laser pulse field. These excitations are analogous to emergent gravitons due to their quadrupole structure and is the excitation of intrinsic metric carried by Hall fluids. Future experiments are needed to reveal how gap structure for fractional Hall systems evolve with finite momentum. Such experiments will not only address the questions related to strongly-interacting electron systems, but can also shed light on the understanding of the analogues of fundamental particles in high energy physics realized in condensed matter systems.
Acknowledgements.— A.K. thanks Avadh Saxena and Pouyan Ghaemi for their fruitful discussions. This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA) under Contract No. 89233218CNA000001. It was supported by LANL LDRD Program and Quantum Science Center, a U.S. DOE Office of Science Quantum Information Science Research Center. It was supported in part by Center for Integrated nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources.