### v3

# Laser-driven Ultrafast Dynamics of a Fractional Quantum Hall System

Abstract
Fractional quantum Hall (FQH) systems are strongly interacting electron systems with topological order. These systems are characterized by novel ground states, fractionally charged and neutral excitations. The neutral excitations are dominated by a low-energy collective magnetoroton mode. Here we derive and use a quasi-one-dimensional model to investigate the ultrafast nonequilibrium dynamics of a laser-driven FQH system within a two-Landau-level approximation. As opposed to the traditional and synthetic bilayers, our model accounts for interactions where electrons can scatter from one Landau-level to another. By performing exact time evolution of the system, we create an out-of-equilibrium state following the laser pulse that shows rich physics. Our calculations show the presence of non-trivial excited modes. One of these modes is electromagnetically active and represent density oscillations of magnetoplasmon mode. Another mode is identified by evaluating the overlap of the initial state and the out-of-equilibrium state following the laser pulse with a quadrupole operator. This mode is analogous to the chiral-graviton mode for FQH systems recently measured in experiments [Nature 628, 78 (2024)]. Our results show that a linearly-polarized pulse field can excite the graviton mode when inter-Landau level scattering occurs.
Introduction.— Fractional quantum Hall (FQH) systems are a prototypical example of strongly correlated physical systems with rich novel phenomena like topological order, quantum geometry and excitations that follow fractional statistics Haldane (2011); Arovas et al. (1984). The neutral excitations of FQH systems are dominated by a collective coherent oscillation mode usually called Girvin-MacDonald-Platzman magnetoroton mode Girvin et al. (1986, 1985). In the long-wavelength limit $k\to 0$, the magnetoroton becomes a quadrupole excitation that carries $L=2$ angular momentum along with the excitation of an underlying quantum geometry metric analogous to the passing of gravitational waves in a medium Haldane (2011). Another novel consequence of the presence of quantum geometry and topology in FQH systems is the quantization of Hall viscosity Haldane (2009); Pu et al. (2020); Read and Rezayi (2011). FQH phase is characterized by the quantization of transverse component of resistivity at low temperatures. This is measured in 2D electrons with density $\rho$ in the presence of high perpendicular magnetic field $B$ at fractional filling $\nu=\rho h/eB$ ($h$ being Planck’s constant and $e$ is the electron charge). Among many odd-denominator fractional filling states observed experimentally Tsui et al. (1982); Stormer et al. (1983) and theoretically explained Laughlin (1983); Haldane (1983), the first observed FQH state at $\nu=\frac{1}{3}$ is of particular interest. It was shown that fractional $\frac{1}{3}$ state can host charged excitations like quasi-particle (quasi-hole) of $e/3$ sometimes called anyons Arovas et al. (1984). It was not until recently that experiments performed on this FQH state gave first direct observation of exchange statistics of fractionally charged particles (anyons) Nakamura et al. (2020). Furthermore, interest in this state is rekindled due to the observation of two $e/3$ quasi-particles scattering into an electron and “Andreev reflected” quasi-hole Glidic et al. (2023). Besides its experimental accessibility, this state has also been readily studied theoretically on different geometries Yoshioka et al. (1983); Rezayi and Haldane (1994); Nakamura et al. (2012); Soulé and Jolicoeur (2012). Moreover, a long wavelength low-energy excitation mode of $\frac{1}{3}$ FQH state has been excited in inelastic linearly-polarized light scattering experiments Pinczuk et al. (1993). This mode was identified later on as a chiral-graviton mode that can be excited through circularly polarized light Liou et al. (2019). Another experiment has also excited this mode through two-photon processes in FQH-1/3 state Liang et al. (2024). Despite such experiments, the theoretical understanding of the dynamics of a multi-Landau-level FQH system has been limited.
In this Letter, we present ultra-fast dynamics of the FQH system is driven out-of-equilibrium by a laser pulse field. The system is initially in $\nu=\frac{1}{3}$ FQH state. We show that when all inter-two-level scattering terms are considered, a short time linearly-polarized light with frequency near Landau level gap can excite chiral-graviton mode below magnetoplasmon mode. We note that while optically driven synthetic bilayers have been recently proposed Ghazaryan et al. (2017); Cian et al. (2020) to host new types of FQH states Halperin (1983); He et al. (1993), as in conventional semiconductor bilayers Eisenstein et al. (1992); Suen et al. (1992); Bishop et al. (2007); Padmanabhan et al. (2009), the electronic systems are assumed to thermalize to the Floquet ground state in the rotating framework of the optical drive field. Therefore, the problem studied in the present work is quite distinct from the situations studied before.
Model Hamiltonian.— We consider a 2D spinless electronic system defined on the surface of torus with lengths $L_{x}$ and $L_{y}$, as shown in Fig. 1 in the presence of both a high perpendicular static magnetic field and a time-dependent electric field from the laser pulse. In the Landau gauge, both magnetic and electric fields can be described by the vector potential $\mathbf{A}=\mathbf{A}_{\rm static}(\mathbf{r})+\mathbf{A}_{\rm EM}(t)$ with $\mathbf{A}_{\rm static}=(0,Bx,0)$ and $\mathbf{A}_{\rm EM}=(0,A_{\rm EM}(t),0)$. The single-particle states in this gauge compatible with torus boundary conditions are given by Yang et al. (2012)
$\phi_{n,j}(\mathbf{r})=\frac{1}{\sqrt{{\cal N}_{n}}}\sum_{k}$ $e^{i\frac{(X_{j}+kL_{x})y}{\ell_{B}^{2}}}e^{-\frac{(x+X_{j}+kL_{x})^{2}}{2\ell_{B}^{2}}}$ (1)
$H_{n}\bigg(\frac{x+X_{j}+kL_{x}}{\ell_{B}}\bigg)$
Where $X_{j}=\frac{2\pi j}{L_{y}}\ell_{B}^{2}$ with $\ell_{B}=\sqrt{\hbar/eB}$ being the magnetic length. ${\cal N}_{n}=L_{y}\sqrt{\pi}2^{n}n!\ell_{B}$ is the normalization constant and $H_{n}$ is Hermite polynomial of order $n$. For each Landau level $n$, index $j$ runs from $0,1,...,N_{\phi}-1$, where $N_{\phi}=L_{x}L_{y}/(2\pi\ell_{B}^{2})$ is the number of flux quanta (Landau orbitals) passing through the system. Within this Landau orbital basis, we are able to write the system Hamiltonian as $\hat{H}=\hat{H}_{sp}+\hat{H}_{int}$, with
$\hat{H}_{sp}$ $=\omega_{c}\sum_{j}\hat{n}_{1,j}+\Gamma_{\rm EM}(t)[\sum_{j}c_{0,j}^{\dagger}c_{1,j}+\mathrm{H.c.}]\;,$ (2)
describing the coupling of electrons with the laser pulse field up to the first Landau level (LL1), and
$\hat{H}_{int}=\sum_{\{n_{i},j_{i}\}}V^{n_{1},n_{2};n_{3},n_{4}}_{j_{1},j_{2};j_{3},j_{4}}c_{n_{1},j_{1}}^{\dagger}c_{n_{2},j_{2}}^{\dagger}c_{n_{3},j_{3}}c_{n_{4},j_{4}}\;,$ (3)
describing the Coulomb interaction with $V_{n_{1},n_{2};n_{2},n_{4}}^{j_{1},j_{2};j_{3},j_{4}}$ the amplitude of two-particle scattering processes due to the repulsive interactions. Here the operators $c_{m,j}$, $c_{m,j}^{\dagger}$ ($\hat{n}_{m,j}{\equiv}c_{m,j}^{\dagger}c_{m,j}$) destroy or create an electron in $m$-th Landau level (LL) at $j$-th orbital localized around $2\pi j\ell_{B}^{2}/L_{y}$ along the axis of the cylinder. The quantity $\omega_{c}$ is the level-spacing between Landau levels 0-1 (see Fig. 1). In each Landau level, cyclotron frequency of electron is quantized in units of $\omega_{c}=eB/m^{*}$ ($m^{*}$ is electron’s effective-mass). The laser field coupling strength $\Gamma_{\rm EM}=\frac{eA_{EM}(t)}{\sqrt{2}m^{*}\ell_{B}}$. We also note for electrons interacting through Coulomb repulsion in the system, the Coulomb energy $E_{c}=e^{2}/\varepsilon\ell_{B}$ ($\varepsilon$ being di-electric constant) is also an important energy scale. The interaction part of our Hamiltonian Eq. (3) contains both intra- and inter-level scattering processes. Throughout the work, we use a short pulse form $A_{\rm EM}(t)=A_{0}e^{-\frac{(t-t_{c})^{2}}{2t_{d}^{2}}}\cos\big[\omega_{0}(t-t_{c})\big]$ where $t_{d}$ is time delay, $t_{c}$ is the center time for the pulse and $\omega_{0}$ is the frequency of incoming laser Fauseweh and Zhu (2020).
Figure 1: (a) Torus with lengths $L_{x}$ and $L_{y}$. Magnetic field is pointing along $\hat{z}$. (b) A schematic of a two-level FQH system. Initially, the system is in fractional $\nu=1/3$ state occupying the LLL ($n=0$). This state is then laser-driven creating excitations to the first Landau level ($n=1$). The energy difference between the two levels is $\omega_{c}$. (c) and (d) are the two-electron scattering processes that do not conserve Landau level index. (c) describes two electrons transitioning from one Landau level to another. (d) a two-electron process where only one electron scatters to another Landau level.
[…]
$V_{j_{1},j_{2},j_{3},j_{4}}^{n_{1},n_{2},n_{3},n_{4}}=\frac{\delta_{j_{1}+j_{2},j_{3}+j_{4}}^{{}^{\prime}}}{2L_{x}L_{y}}\sum_{\mathbf{q},\mathbf{q}\neq 0}\delta_{q_{y},t\frac{2\pi}{b}}\delta_{q_{x},s\frac{2\pi}{a}}\delta_{j_{1}-j_{4},t}^{{}^{\prime}}\frac{2\pi e^{2}}{\varepsilon q}C_{n_{1},n_{4}}(-\mathbf{q})C_{n_{2},n_{3}}(\mathbf{q})e^{-i2\pi s\frac{j_{1}-j_{3}}{N_{\phi}}}e^{-\frac{q_{x}^{2}+q_{y}^{2}}{2}\ell_{B}^{2}}$ (4)
[…]
Dynamical observables. Our system is initially in FQH-$\frac{1}{3}$ state $|\Psi_{0}\rangle$ at time $t=0$. We then turn on the laser drive and let the system evolve through Schrodinger equation as $|\Psi(t)\rangle=\hat{\mathit{T}}e^{-i\int_{0}^{t}\hat{H}(t^{\prime})dt^{\prime}}|\Psi_{0}\rangle$, where $\hat{\mathit{T}}$ is the time ordering operator and $|\Psi(t)\rangle$ is our out-of-equilibrium state. The time-ordering is achieved by discretizing the time-dependent matrix exponential in small time steps. The dynamical response of the system can be understood with time-dependent structure factor ${\mathcal{S}}(\mathbf{q})=\langle\hat{\rho}_{\mathbf{q}}\hat{\rho}_{-\mathbf{q}}\rangle$ where $\rho_{\mathbf{q}}$ is the projected density (see SM  sup ).
Results. We consider level-spacing $\omega_{c}<E_{c}$, i.e. the level spacing is less than Coulomb energy scale. The ratio of these energies $\kappa\equiv\frac{e^{2}}{\hbar\omega_{c}\epsilon\ell_{B}}$ is magnetic field dependent ranging from $2.6/\sqrt{B[T]}$ for Ga-As to $22.5/\sqrt{B[T]}$ for Al-As justifying the our approximation for fractional $\nu=1/3$ Hall system usually observed at $B\sim 10$ Tesla Sodemann and MacDonald (2013).
Figure 2: Color online. Structure factor in frequency domain ${\mathcal{S}}_{\mathbf{q}}(\omega)$ for different system sizes. The torus dimensions is taken $L_{x}=L_{y}=\sqrt{6\pi N}\ell_{B}$. The laser-frequency is $\omega_{0}=0.7\omega_{c}$. The intensity is set to $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/E_{c}$. The units of time is the inverse of $E_{c}$. The usual values of $E_{c}$ is around $14\;{\rm meV}$ Pinczuk et al. (1993) which gives 4747 femto-seconds as the unit of time. The structure factor spectrum is normalized such that $\int{\mathcal{S}}_{\mathbf{q}}(\omega)d\omega=1$.
The spectrum of structure factor ${\mathcal{S}}_{\mathbf{q}}(\omega)$ is shown in Fig. 2. The peak at the level-spacing (cyclotron-energy) $\omega_{c}$ is the plasmon equivalent of FQH state identified as $q\to 0$ limit of magnetoplasmon mode (see SM sup for other quantities). Significantly, there emerge new modes located at the energy around $0.2\omega_{c}$ for different system sizes. To determine the nature of theses new low-energy modes in $S_{\mathbf{q}}(\omega)$, we evaluate the overlap $|\langle{\Psi(t)}|n\rangle|$ to analyze the weight of low-lying eigen-states $|n\rangle$ contained in the photo-induced excited state $|\Psi(t)\rangle$ as shown in Table 1. We have found that low-energy excitations are dominated by quadrupole FQH graviton states. These states are identified by matrix-element $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$, where $\hat{O}^{(2)}$ is the operator with quadrupole structure and $|\Psi_{0}\rangle$ is our initial ground state before the application of laser pulse. The specific form of $\hat{O}^{(2)}$ is given as Liou et al. (2019)
$\hat{O}^{(2)}=\sum_{\{j_{i}\}}\delta_{j_{1}+j_{2},j_{3}+j_{4}}^{{}^{\prime}}\hat{O}^{j_{1},j_{2},j_{3},j_{4}}_{qd}c_{0,j_{1}}^{\dagger}c_{0,j_{2}}^{\dagger}c_{0,j_{3}}c_{0,j_{4}}\quad$ (5)
$\hat{O}^{\{j_{i}\}}_{qd}=\sum_{q=(s\frac{2\pi}{L_{x}},t\frac{2\pi}{L_{y}})}^{{}^{\prime}}\delta_{j_{1}-j_{4},t}^{{}^{\prime}}\bigg(q_{x}^{2}-q_{y}^{2}\bigg)\frac{2\pi}{q}e^{-\frac{q^{2}}{2}}e^{-i2\pi s\frac{j_{1}-j_{3}}{N_{\phi}}}$
Figure 3: Color online. (a) $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$ for $N=7$ for $L_{y}=\sqrt{2\pi 21}\ell_{B}$. Where $|\Psi_{0}\rangle$ is the FQH-1/3 state and $|n\rangle$ is the $n$-th excited state of Hamiltonian in Eq. (3). The inset of (a) gives the transition matrix element between FQH-1/3 state and the eigen-states of LLL. (b) Same as (a) but for $N=6$ and $L_{y}=\sqrt{2\pi 18}\ell_{B}$. The level-spacing is set to $0.5E_{c}$. Grey vertical line on plots is at energy $0.14E_{c}$.
Table 1: The contribution of states in the photo-induced state $|\Psi(t)\rangle$ after the pulse for N=N=5, 6, 7 and 8. $\Delta E$ denotes the energy difference from the ground state. Overlap $|\langle n|\Psi(t)\rangle|$ gives the contribution of individual eigenstate $|n\rangle$ of the system Hamiltonian with energy $E_{n}$. $\omega_{c}=0.5E_{c}$.
Size  [^1] The parameters of the laser pulse are set to $(eA_{0}/\sqrt{2}m^{*}\ell_{B},\omega_{0})=(0.2E_{c},0.7\omega_{c})$ with $t_{d}=10E_{c}$ and $L_{y}=\sqrt{2\pi N_{\phi}}\ell_{B}$. Similar results have been found for observed for near-TT limit. $E_{n}$ ($E_{c}$) Overlap $\Delta E$ $(E_{c})$ Nature
N=5 -1.067 0.883 0.000 G.S.
- -0.938 0.097 0.129 Quadrupole
- -0.852 0.01 0.215 Quadrupole
- -0.562 0.4 $0.504\approx\omega_{c}$ Plasmon
Total weight 0.097 Quadrupole
$N=6$ -1.381 0.850 0.000 G.S.
- -1.296 0.095 0.085 Quadrupole
- -1.283 0.041 0.098 Quadrupole
- -1.235 0.028 0.146 Quadrupole
- -1.210 0.034 0.171 Quadrupole
- -1.176 0.014 0.205 Quadrupole
- -0.876 0.429 $w_{c}$ Plasmon
Total weight 0.212 Quadrupole
N=7  [^2] Due to large Hilbert space, for $N=7$, we only have low-energy quadrupole states by which we identify graviton modes in the structure factor. For $N=8$, we have only have eigenvectors corresponding to first 100 lowest energy eigenstates that include few graviton states. It is expected that for $N=8$ States that give more than $0.01$ overlap with $|\psi(t)\rangle$ are considered. -1.702 0.827 0.000 G.S.
- -1.606 0.058 0.096 Quadrupole
- -1.587 0.065 0.115 Quadrupole
- -1.583 0.046 0.119 Quadrupole
- -1.562 0.013 0.140 Quadrupole
- -1.548 0.030 0.161 Quadrupole
- -1.530 0.024 0.172 Quadrupole
- -1.2 - $w_{c}$ Plasmon
Total weight 0.236 Quadrupole
N=8  [^2] Due to large Hilbert space, for $N=7$, we only have low-energy quadrupole states by which we identify graviton modes in the structure factor. For $N=8$, we have only have eigenvectors corresponding to first 100 lowest energy eigenstates that include few graviton states. It is expected that for $N=8$ States that give more than $0.01$ overlap with $|\psi(t)\rangle$ are considered. -2.035 0.797 0.000 G.S.
- -1.951 0.026 0.084 Quadrupole
- -1.926 0.049 0.109 Quadrupole
- -1.909 0.063 0.126 Quadrupole
- -1.896 0.013 0.139 Quadrupole
- -1.884 0.038 0.151 Quadrupole
- -1.873 0.019 0.162 Quadrupole
Total weight 0.208 Quadrupole
In Fig. 3, we present the values of transition matrix element $\langle n|\hat{O}^{(2)}|\Psi_{0}\rangle$ (un-normalized) between FQH 1/3 state and $n$-th excited state $|n\rangle$ for two cases i) When $|n\rangle$ is the eigen-state of the system Hamiltonian involving all scattering processes and ii) when $|n\rangle$ is the eigen-state of LLL Hamiltonian where all scattering processes are confined to LLL. The transition matrix element $I_{n}$ represents the system transition rate due to an oscillating metric tensor analogous to the passing of a gravitation wave and hence usually termed as graviton Yang (2016). Fig.  3 (a) and (b) shows the presence of such mode(s) for various system sizes. This mode occurs for both liquid limit and near-TT limit (See SM sup ). As seen from Fig. 3, the energy of graviton states $E_{grav}$, when all scattering processes are included, is slightly lower than the graviton states of a system when all interaction processes are confined to LLL. Furthermore, we follow Ref. Liou et al., 2019 to analyze the handedness of this graviton and find that it has the structure of $(q_{x}-iq_{y})^{2}$, suggesting its chiral nature (See SM  sup for the detail of analysis.). Previously, it has also been proposed  Kirmani et al. (2022); Liu et al. (2018) that this low-energy mode can be excited via geometric quench Liu et al. (2021). A variant of such graviton mode called chiral graviton mode has been observed in experiments recently for fractional Hall states at filling fractional $\frac{1}{3}$ and $\frac{2}{5}$ fillings Liang et al. (2024). Chiral graviton modes in FQH system has been previously excited by tuning the laser-frequency near resonance conditions between valence and conduction band in 2D quantum wells. Our work has demonstrated that chiral graviton modes can also be exited when the a short-time laser is tuned near Landau level gap which is easily accessible experimentally. Furthermore, the excitation of graviton mode can be observed along with magneto-plasmon mode when all scattering processes are considered as seen from the structure factor.
Figure 4: The structure factor $S_{\mathbf{q}}(\omega)$ (unnormalized) for $\mathbf{q}=(2\pi/L_{x},0)$, $\omega=0.7\omega_{c}$ and $N=6$ electron system. All parameters are set according to the Fig. (2). Without the two-particle chiral excitations (Fig. 1 (c)) the intra-$L$ mode is suppressed. Inset gives the strength of the main peak of the low-energy (graviton) and plasmon mode against the intensity of the laser-pulse. Quadratic dependence and linear dependence is evident for the low energy mode and plasmon mode respectively. The peak value of the plasmon mode is divided by 5 for better fit.
The appearent drop in the dynamical structure factor when normalized in Fig. 2 between $N=6$ and $N=7$ is due to the presence of more quadrupole excitations present in $N=7$ photo-excited state $|\Psi(t)\rangle$. As seen from Fig. 3 (a) and (b), there are more graviton states for $N=7$ consistent with Ref. Liou et al., 2019. These states are also identified in Table. 1. Most importantly, the total low-energy quadrupole overlap in the photo-induced state does not drop with the system size as seen from Table. 1. We have shown that the quadrupole excitations occur up-to $N=8$.
The two-particle process that excites the low-energy mode(s) through photo excitation, is given in Fig. 1 (c). This process describes inter-level two-electron transitions from one Landau level into another thus requiring two-photons. Moreover, this process has amplitude that is chiral in nature $\frac{(q_{x}-iq_{y})^{2}}{2\sqrt{q_{x}^{2}+q_{y}^{2}}}e^{-\frac{q^{2}}{2}}c_{1,j_{1}}^{\dagger}c_{1,j_{2}}^{\dagger}c_{0,j_{3}}c_{0,j_{4}}+\mathrm{H.c.}$ (See SM sup ). As seen in Fig. 4, when such (chiral) two-particle excitations are omitted, there is a significant drop in the low-energy mode present in the structure factor. The dependence of the low-energy and the plasmon peak strength with the laser intensity in the limit $A_{0}\to 0$ is also shown in inset of Fig. 4. It is evident that in the low intensity regime, the magnetoplasmon and the low-energy mode follow the linear and quadratic dependence of the field strength respectively. This is in accordance with the fact that graviton couple with two-particle processes while magnetoplasmon depends upon single-photon process. Quantifying such quadratic dependence gives an alternate avenue to measure the graviton modes through non-linear absorption experiments. One can then use a circularly polarized probe light to experimentally detect the chirality of the excited graviton modes if needed.
Conclusions.— Through exact-diagonalization, the laser-induced dynamics of FQH state at $\nu=1/3$ has been presented. We showed that laser-induced excitations of FQH states have rich physics. On the one hand, some excitations are coupled to electromagnetic field and can cause emission like the plasmon mode. On the other hand, we have shown that when scattering from LLL to LL1 is included in interactions, it is possible to excite neutral-modes of LLL through linearly polarized laser pulse field. These excitations are analogous to emergent gravitons due to their quadrupole structure and is the excitation of intrinsic metric carried by Hall fluids. Moreover, we have shown that non-linear light absorption experiments are an alternate route to the detection of such modes being quadratically dependent upon the light’s intensity. Future non-linear absorption experiments are needed to probe such excitations. This can also help with understanding how the gap structure for fractional Hall systems evolve with finite momentum. Such experiments will not only address the questions related to strongly-interacting electron systems, but can also shed light on the understanding of the analogues of fundamental particles in high energy physics realized in condensed matter systems.
Acknowledgements.— This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA) under Contract No. 89233218CNA000001. It was supported by Quantum Science Center, a U.S. DOE Office of Science Quantum Information Science Research Center, and in part by Center for Integrated nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. A.K. thanks Avadh Saxena and Pouyan Ghaemi for their fruitful discussions. A.K. was supported LANL LDRD Program.